find the slope of the tangent line to f(x)=x2 + 2x at x=3


1
Expert's answer
2021-01-28T05:35:15-0500

Solution:

Given: "f(x)=x^2+2x" ...(i)

Slope of tangent is "f'(x)"

So, we find derivative of given "f(x)"

On differentiating (i) w.r.t "x"

"f'(x)=2x+2"

Putting "x=3" in "f'(x)" , we get,

"f'(3)=2(3)+2=6+2=8"

Answer: 8